OpenStudy (anonymous):

The emergent velocity, v, of a liquid flowing from a hole in the bottom of a tank is given by v = Square root(2gh) where g is the acceleration due to gravity and h is the depth of the liquid in the tank. Find the rate of change of v with respect to h when (a) h = 9 , (b) h = 4 .

OpenStudy (mathmate):

\( v=\sqrt{2gh} \) can you find \( \large \frac{dv}{dh} \) ?

OpenStudy (anonymous):

Would it be \[\frac{ 1 }{ 2 }(2gh)^{-\frac{ 1 }{ 2 }}\] ?

OpenStudy (mathmate):

You can take the constant \( \sqrt{2g} \) out before you integrate.

OpenStudy (mathmate):

*differentiate.

OpenStudy (anonymous):

And have it as \[\sqrt{2g} + \sqrt{h} \] ?

OpenStudy (mathmate):

"Take out" is meant to factorize and put it on the left of the differentiation: \( \frac{d\sqrt{2gh}}{dh} \\ =\sqrt{2g}\frac{d\sqrt{h}}{dh} \\ =\sqrt{2g} {\frac{1}{2}}\frac{1}{\sqrt{h}} \\ =\frac{\sqrt{g}}{2h} \)

OpenStudy (anonymous):

Ah, that is far more elegant. lol

OpenStudy (anonymous):

Do I just input the values now?

OpenStudy (mathmate):

No, there is a typo. it should read \( \frac{\sqrt{2g}}{2\sqrt{h}} \) Typo is all on the last line.

OpenStudy (anonymous):

That makes much more sense.

OpenStudy (mathmate):

Do check answers and be convinced that they are correct before you put it into your answer, i.e. before you "own" it.

OpenStudy (anonymous):

I usually do. I was sure I was missing something and have been thinking it though the entire time. I thought you had pulled out the square root of two and somehow turned it into two on the bottom. It made my brain fart. ><;

OpenStudy (mathmate):

That's good. Getting the right answer is a temporary pleasure. Knowing how to solve the problem is the real success! Keep up the good work!

OpenStudy (anonymous):

So, since that is the first derivative, that represent rate of change, so plug in the force of gravity (32 ft/sec) and the two h's given, right? (Yeah! Or in Isaac Asimov's words: "The true pleasure is in the finding out, rather than the knowing." :D

OpenStudy (mathmate):

Exactly! Note: are you working in imperial system or metric? 32.2 ft/s^2 or 9.81 m/s^2 ?

OpenStudy (mathmate):

I agree with that too!

OpenStudy (anonymous):

Imperial. Mayhaps someday the United States will convert to metric. >.>